OpenStudy (anonymous):

adding fractions this is my math problem. 10/15 + 2/15

OpenStudy (radar):

If they all have the same denominator you may combine them appropriately with their sign. In the case above they are common, both being 15 you can just add getting 12/15 You can divide top and bottom without changing the value getting 5/5 which is equivalent to 12/15 or is the reduced version.

OpenStudy (radar):

You shouldn't get 5/5 when you divide by 3 how about 4/5 lol

OpenStudy (anonymous):

ok. how about 1/3 +1/6=

OpenStudy (anonymous):

this problem 2/11 + 7/11 = 9/11

OpenStudy (radar):

1/3 + 1/6 Needs a common denominator. It makes sense to get one that is small as possible to keep the problem simple. I am looking for a number that is as small as possible that has a factor 6 and 3. This one is easy as the 6 will contain the 6 and it also contains the 3 (3 x 2 = 6)

OpenStudy (radar):

Yes the 9/11 was correct. To the problem at hand, what do you need to multply the denominator 3 to get 6? I'll be back later

OpenStudy (radar):

If you reasoned it to be 2 you are correct. Multiply both top and bottom by 2 getting 2/6 since the second fraction already has 6 we can play we multiplied top and bottom by 1. 2/6 +1/6 = 3/6 we can reduce this by dividing top and bottom by 3 getting 1/2